OpenStudy (anonymous):

Solve the quadratic equation 12x2 − 3x − 9 = 0 by factoring

OpenStudy (theviper):

\[\LARGE{12x^2-3x-9=0}\]\[\implies 12x^2-12x+9x-9=0\]

OpenStudy (theviper):

\[12x(x-1)+9(x-1)=0\]

OpenStudy (theviper):

\[(12x+9)(x-1)=0\]

OpenStudy (theviper):

\[12x+9=\frac{0}{x-1}=0\]

OpenStudy (theviper):

\[x=\frac{-9}{12}\]

OpenStudy (theviper):

& other x=1

OpenStudy (theviper):

got it??? :)

OpenStudy (theviper):

????